How many milliliters of 3 F KOH can be prepared from 100 grams of KOH?


1
Expert's answer
2021-03-03T01:32:41-0500

Formality is expressed as:

F = n / V = m / (Mr × V)

where F - formality, n - number of moles, m - mass, Mr - molecular weight, V - volume

From here:

V = m / (Mr × F)

V(NaOH) = m(NaOH) / [Mr(NaOH) × F(NaOH)] = 100 g / (40 g/mol × 3 F) = 0.83 L = 83 mL


Answer: 83 mL